# Distilling Reasoning Ability from Large Language Models with Adaptive Thinking

### v5

@@ IV-A1 Datasets
12 complex reasoning datasets divided into four classes are used to evaluate PST: arithmetic (SingleEq [53], AddSub [54], MultiArith [55], GSM8K [56], Aqua[57], Svamp [58]), symbolic [25] (Last Letter Concatenation, Coin Flip), common sense (CommonSenseQA [59], StrategyQA [60]) and other (Date Understanding, Track Shuffled Objects). The train-test splitting of these datasets follows ho et al.[6]. For the sake of simplicity, we use abbreviations (concatenated initials) to represent the name of each dataset in Table I and Table II.
@@ IV-A3 Training Details
The student model is implemented by PyTorch [62] and trained on 2 NVIDIA 4090 GPUs. When training model, the batch size, initial learning rate and training epoch are set to 4, 1e-5 and 20 respectively. For varying learning rate, we apply cosine with restarts whose warm-up step is 1200. For optimizing model parameters, adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$ and $weight\_decay=0.1$ is adopted. And the max rationale length is limited to 300. As for the $K$, we set it to 50 across all tasks due to the better performance of this setting.
@@ IV-A4 Evaluation Metric:
to evaluate the performance of different methods, where $TP$, $FP$, $FN$ and $TN$ indicate true positive, false positive, false negative and true negative counts respectively.
@@ IV-B2 Effectiveness of submodules in ATM
TABLE IV: The performance of different strategies on GPT2-Large and T5-Large. Accuracy (%) is reported across 4 types of representative reasoning tasks.
@@
TABLE V: The display about how the model (GPT-Large) trained with ATM to divide the thinking mode on the StrategyQA test set. 10 pre-thinking and 10 post-thinking questions were randomly selected, respectively.
Type Questions
Pre-Thinking 1. In a hypothetical race between a Swallow and an American Woodcock, would the Swallow win?
@@ 10. Would you be likely to see storks at a baby shower?
Post-Thinking 1. Is Christmas celebrated during winter?
@@ IV-C2 Attention map
second, we visualize the attention map between the token that is used to generate answers and other tokens in Fig. 4. Through the attention map, one can first find that the answer generation overly relies on the rationale under the pre-thinking setting because almost all tokens with higher attention values are included in the rationale, which causes the correctness of the answer to be seriously affected even though just minor errors occur in rationale. In addition, the tokens with higher attention value in the question are tend to occur at the end of the question as shown in the position corresponding to the red vertical dotted line of the Fig. 4, which indicates that pre-thinking compresses the question information to the last token of the question so that the question information that the SLM can notice when generating the answer is lossy. Compared with pre-thinking, post-thinking can focus more on the complete information of the question itself when generating answers, and can escape from the direct influence of minor errors in rationale. Furthermore, it can be observed that the attention sparsity in the attention map of from post-thinking to prefixes mechanism increases, indicating that pre-thinking can notice more key points in the question.
@@ IV-C3 Efficiency
thirdly, we compare the inference speed between pre-thinking and post-thinking on three different datasets in Fig.5. The inference speed of post-thinking is at least 50 times faster than that of pre-thinking, which is important for practical tasks that are sensitive to model inference time. In addition, since the average lengths of rationales generated for different types of questions are diverse, the average time spent for pre-thinking on different datasets also varies greatly.
Figure 5: Inference speed of GPT2-Large on 3 datasets with different cot-finetuning strategies. As shown in the figure, since the post-thinking can stop inference after giving a precise answer, its speed is much faster than the compared pre-thinking method.
@@ IV-C4 Case study
In order to further evaluate the perception ability of ATM on the complexity of input questions, 10 questions are randomly selected from each of the two question sets used for pre-thinking and post-thinking respectively in ATM, and are shown in Table V. In the table, questions answered through pre-thinking are typically of higher complexity, often requiring SLM to decompose the questions, engage in reasoning, and perform comparative analyses to obtain answers. Conversely, questions answered using post-thinking in the table are of lower complexity and can often be directly resolved. Thus, this adaptive mechanism in ATM, which selects thinking approaches based on question complexity, is considered to be more flexible and closely aligned with human question-answering methods compared to pre-thinking and post-thinking.
However, in Table V, from a human perspective, one can also observe that the set of questions using post-thinking contains several questions (7, 9) that appear complex, while the other question set includes some questions (6, 8, 10) with lower complexity. This is related to the model’s inherent question-solving ability. Influenced by factors such as training data and methods during the pre-training, different SLM possess varying foundational capabilities. Since ATM essentially categorize the question sets based on the SLM’s own abilities, it leads to minor differences in question categorization between ATM and human perception of question complexity.
@@ IV-E Comparison with Other Methods
We compare against other state-of-the-art models related to us on 4 types of datasets in Table IV. Below are a brief introduction to these models.
Distilling Step-by-Step [7]: it prepends a prefix ([label] or [rationale]) to the question to distinguish whether the current task is to output an answer or generate a rationale.
Mint [64]: Mint is a multi-task model based on Distilling Step-by-Step, specifically designed for arithmetic reasoning questions. In addition to rationales in natural language form, the SLM is also trained to output rationales in the form of equation solutions and solution tree pre-order traversal. ‘-’ in Table IV means that Mint cannot be applied to tasks other than arithmetic reasoning.
Fine-tune-CoT [6]: for a single question, it allows the teacher LLM to generate multiple rationales to train the student SLM, so that making SLM can understand the core reasoning logic contained in the rationale. The number of rationales for a single question is 8 in our experiment.
SCOTT [41]: by introducing the counterfactual training samples, SCOTT makes the final answer produced by SLM faithful to the intermediate rationale. Note that the contrastive decoding strategy proposed by SCOTT was discarded in our experiment because the parameters of gpt-3.5-turbo are not open-source.
Specializing [39]: Specializing improves the performance of SLM by adjusting the training samples into different data forms including zero-shot answer-only, in-context answer-only, in-context chain of thought, and zero-shot chain-of-thought.
Furthermore, in the table, w. post-thinking refers to that transferring training object related to Equation 1 in above methods to 2, and w. ATM represents that the strategies proposed by the above methods are used in steps 2, 3, and 8 of Algorithms 1, respectively.
@@
In the table, ATM shows comparable performance to existing excellent cot-finetuning methods. Second, The plug-and-play improvement of pre-thinking and ATM over these methods is also demonstrated.

### v6

@@ IV-A1 Datasets
12 complex reasoning datasets divided into four classes are used to evaluate the proposed method: arithmetic (SingleEq [63], AddSub [64], MultiArith [65], GSM8K [66], Aqua[67], Svamp [68]), symbolic [69] (Last Letter Concatenation, Coin Flip), common sense (CommonSenseQA [70], StrategyQA [71]) and other (Date Understanding, Track Shuffled Objects). The train-test splitting of these datasets follows ho et al.[6]. For the sake of simplicity, we use abbreviations (concatenated initials) to represent the name of each dataset in Table I, Table II, and Table X.
@@ IV-A3 Training Details
The student model is implemented by PyTorch [73] and trained on 2 NVIDIA 4090 GPUs. When training model, the batch size, initial learning rate and training epoch are set to 4, 1e-5 and 20 respectively. For varying learning rate, we apply cosine with restarts whose warm-up step is 1200. For optimizing model parameters, adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$ and weight_decay=0.1weight\_decay=0.1 is adopted. And the max rationale length is limited to 300. As for the $K$, we set it to 50 across all tasks due to the better performance of this setting.
@@ IV-A4 Evaluation Metric:
to evaluate the performance of different methods, where $TP$, FPFP, FNFN and $TN$ indicate true positive, false positive, false negative and true negative counts respectively.
@@ IV-B2 Effectiveness of submodules in ATM
IV-B3 Effectiveness of ATM on more advanced SLMs
To explore whether the proposed method is also effective on more advanced SLMs, we additionally conduct experiments on two models: Qwen2.5-0.5B[75] and LLama3.2-1B[76]. As shown in Table IV, ATM improves performance on both models, demonstrating the generality and robustness of our approach.
TABLE IV: The performance of different strategies on Qwen2.5-0.5B and LLama3.2-1B. Accuracy (%) is reported on StrategyQA.
Methods Pre Post Qwen2.5-0.5B LLama3.2-1B
Standard Finetuning 64.77 64.91
Pre-thinking $\checkmark$ 64.91 66.08
Post-thinking $\checkmark$ 65.35 67.24
ATM $\checkmark$ $\checkmark$ 66.08 67.78
TABLE V: The performance of different strategies on GPT2-Large and T5-Large. Accuracy (%) is reported across 4 types of representative reasoning tasks.
@@ IV-B3 Effectiveness of ATM on more advanced SLMs
Prompt Engineering CoT[5] 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
CoT w. gpt-3.5-turbo 88.33 68.44 64.66 53.71
COD[77] 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
COD w. gpt-3.5-turbo 82.22 28.88 37.33 54.14
SC[30] 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
SC w. gpt-3.5-turbo 91.11 72.44 68.00 55.74
Self-Refine[78] 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Self-Refine w. gpt-3.5-turbo 81.66 - - -
@@
TABLE VI: The efficiency and complexity of different strategies on GPT2-Large. Parameters (M), reasoning time (S/Sample), average accuracy (%), training GFLOPs, and training time (Hours) are reported across 4 types of representative reasoning tasks (MA, TSO, LLC, and SQA). Vanilla post-thinking and vanilla ATM mean that no additional strategies are plugged in, while other post-thinking and ATM methods are combined with corresponding strategies.
Mechanism Methods Parameters Reasoning Time Accuracy GFLOPs Training Time
Pre-thinking Fine-tune-CoT [6] 774 9.8 33.07 784.09 7
SCOTT [43] 774 8.9 33.75 784.09 4
Specializing [39] 774 8.8 32.35 993.44 7
Prefixes Distilling Step-by-Step [7] 774 0.1 32.84 1397.8 3
Mind’s Mirror [36] 774 0.1 35.23 1397.8 3
Post-thinking Vanilla 774 0.1 33.61 784.09 2
Fine-tune-CoT [6] 774 0.1 34.02 784.09 7
SCOTT [43] 774 0.1 35.11 784.09 4
Specializing[39] 774 0.1 33.47 993.44 7
ATM Vanilla 778 6.1 34.70 784.09 6
Fine-tune-CoT [6] 778 6.1 35.65 784.09 21
SCOTT [43] 778 6.0 37.08 784.09 12
Specializing [39] 778 6.2 34.76 993.44 22
TABLE VII: The display about how the model (GPT-Large) trained with ATM to divide the thinking mode on the StrategyQA test set. 10 pre-thinking and 10 post-thinking questions were randomly selected, respectively.
Pre-thinking Questions
1. In a hypothetical race between a Swallow and an American Woodcock, would the Swallow win?
@@ 10. Would you be likely to see storks at a baby shower?
Post-thinking Questions
1. Is Christmas celebrated during winter?
@@ IV-C2 Attention map
Second, we visualize the attention map between the token that is used to generate answers and other tokens in Fig. 4. Through the attention map, one can first find that the answer generation overly relies on the rationale under the pre-thinking setting because almost all tokens with higher attention values are included in the rationale, which causes the correctness of the answer to be seriously affected even though just minor errors occur in rationale. In addition, the tokens with higher attention value in the question are tend to occur at the end of the question as shown in the position corresponding to the red vertical dotted line of the Fig. 4, which indicates that pre-thinking compresses the question information to the last token of the question so that the question information that the SLM can notice when generating the answer is lossy. Compared with pre-thinking, post-thinking can focus more on the complete information of the question itself when generating answers, and can escape from the direct influence of minor errors in rationale. Furthermore, it can be observed that the attention sparsity in the attention map from post-thinking to prefixes mechanism increases, indicating that post-thinking can notice more key points in the question when compared with prefixes mechanism.
@@ IV-C3 Efficiency
Thirdly, we compare the inference speed between pre-thinking and post-thinking on three different datasets in Table VI. The inference speed of post-thinking is at least 50 times faster than that of pre-thinking, which is important for practical tasks that are sensitive to model inference time.
@@ IV-C4 Case study
In order to further evaluate the perception ability of ATM on the complexity of input questions, 10 questions are randomly selected from each of the two question sets used for pre-thinking and post-thinking respectively in ATM, and are shown in Table VII. In the table, questions answered through pre-thinking are typically of higher complexity, often requiring SLM to decompose the questions, engage in reasoning, and perform comparative analyses to obtain answers. Conversely, questions answered using post-thinking in the table are of lower complexity and can often be directly resolved. Thus, this adaptive mechanism in ATM, which selects thinking approaches based on question complexity, is considered to be more flexible and closely aligned with human question-answering methods compared to pre-thinking and post-thinking.
However, in Table VII, from a human perspective, one can also observe that the set of questions using post-thinking contains several questions (7, 9) that appear complex, while the other question set includes some questions (6, 8, 10) with lower complexity. This is related to the model’s inherent question-solving ability. Influenced by factors such as training data and methods during the pre-training, different SLM possess varying foundational capabilities. Since ATM essentially categorize the question sets based on the SLM’s own abilities, it leads to minor differences in question categorization between ATM and human perception of question complexity.
@@ IV-E Comparison with Other Methods
We compare the performance and efficiency of our approach with 3 other types of state-of-the-art models related to ours on 4 datasets in Tables V and VI. The first type of comparison method is the prefixes mechanism including distilling Step-by-Step [7], Mint [37], MMI [38], and Mind’s Mirror [36]. ‘-’ of Mint and MMI in Table V means that Mint cannot be applied to tasks other than arithmetic reasoning and MMI releases their implementation only based on the t5 framework, respectively. The second type of comparison method is the pre-thinking mechanism including Fine-tune-CoT [6], SCOTT [43] and Specializing [39]. We integrate all the above methods into our unified training framework and data based on their open source implementations for fair comparison. Another type of comparison method is prompt engineering including CoT [5], CoD [77], SC [30], and Self-Refine [78]. All methods of prompt engineering are tested on both LLM(w. gpt-3.5-turbo in Table V) and SLM. Furthermore, in the table, w. post-thinking refers to that transferring training object related to Equation 1 in pre-thinking methods to 2, and w. ATM represents that the strategies proposed by the above pre-thinking methods are used in steps 2, 3, and 8 of Algorithms 1, respectively.
IV-E1 Accuracy
@@
In Table V, one can first observe that although prompt engineering enables LLM to achieve good performance on different reasoning tasks, they do not work on SLM, which proves the necessity of CoT-distillation. Second, ATM shows comparable performance to existing excellent distillation methods. Moreover, The plug-and-play improvement of post-thinking and ATM over pre-thinking methods is also demonstrated.
IV-E2 Efficiency and Complexity
In Table VI, a comprehensive comparison of the efficiency and complexity of ATM and other methods are displayed. For the training phase, the perception module in ATM results in only a limited increase in overall space (parameters) and time (GFLOPs) complexity. And the overall training time increase as ATM requires training two additional models for the pre-thinking and post-thinking modes during data partitioning. In summary, although ATM incorporates additional components and strategies during training, the overall overhead remains within a reasonable range. As for the practical deployment of SLM, memory consumption and inference speed during testing are more critical. As shown in Table VI, ATM slightly increases memory usage due to the additional parameters introduced in the perception module. In terms of inference speed, ATM benefits from the post-thinking mechanism, making its overall inference speed significantly faster than pre-thinking-based methods. Among all methods, prefixes and post-thinking achieve the fastest inference speed. However, their overall performance is lower than ATM, and prefixes suffers from inconsistencies between the reasoning process and the final answer. In conclusion, ATM maintains the low-resource requirements of SLM while achieving the best trade-off between speed and accuracy.
Furthermore, we also compare ATM with other representative human-inspired dual-system thinking framework, comprising fast and slow thinking. We apply their official implementations to reproduce the result. As shown in Table VIII, ATM outperforms these baselines, thereby further validating its effectiveness and advantages.
TABLE VIII: Comparison of different human-inspired dual-system thinking frameworks. The student model is GPT2-Large and the accuracy (%) of each dataset is reported.
Method MA TSO LLC SQA
Dualformer[52] 23.88 43.11 4.00 64.19
LS-Mixture[48] 21.66 41.33 4.00 63.75
ATM 25.55 43.55 4.66 65.06
IV-F Discussion
TABLE IX: A comparison of the quality of rationales generated by different teacher LLMs and their impact on the student SLM GPT2-Large.
Teacher Intermediate Errors(%) Other Errors (%) Accuracy of SLM (%)
gpt-3.5-turbo 40 4 43.55
Multi agent debate[79] 20 0 49.78
TABLE X: Comparison of different distillation mechanisms on the unified cross-lingual and cross-task reasoning dataset. The overall accuracy (%) and the accuracy (%) of each subset of datasets are reported.
Method MA TSO LLC SQA CMATH AVERAGE
Standard Finetuning 25.00 37.77 7.33 63.17 3.88 27.43
Pre-thinking 33.33 46.22 21.33 60.26 3.33 32.89
Post-thinking 27.22 43.11 15.33 64.19 3.88 30.74
ATM 35.55 49.77 22.66 65.21 4.44 35.52
TABLE XI: Accuracy (%) and GFLOPs on MA with different features fed into perception module. The student model is GPT2-Large. $\checkmark$ represents that the corresponding feature is used.
Method Query Tokens Question Tokens Word Count Readability MDD Accuracy GFLOPs
ATM $\checkmark$ 23.88 790.65
ATM $\checkmark$ $\checkmark$ 25.00 791.38
ATM $\checkmark$ $\checkmark$ $\checkmark$ 25.55 791.43
ATM $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 25.55 791.44
ATM $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 26.11 791.48
Figure 6: Case study on intermediate reasoning errors. Text highlighted in yellow indicates incorrect reasoning steps, while text highlighted in green represents correct intermediate steps. The green answer indicates that the answer is correct. The blue, red, and green robot icons correspond to the teacher LLM (gpt-3.5-turbo), the SLM trained on rationales generated by gpt-3.5-turbo, and the SLM trained on rationales generated through the multi-agent debate [79], respectively.
IV-F1 Better reasoners are better teachers
Although ATM incorporates answer information into a prompt to enhance the quality of the rationale generated by the teacher LLM, biased or suboptimal rationales are still inevitable. As shown in the first row of Table IX, we conduct a quantitative analysis on 50 randomly sampled cases from the TSO dataset. The results indicate that a primary issue in the LLM-generated rationale is minor errors in intermediate reasoning steps. Fig. 6 demonstrates this bias inherited by the student SLM from the teacher LLM.
To mitigate this issues, integrating multiple teacher LLMs as a better reasoner is a feasible solution to reduce bias in the rationale. For an instance, we employ multi agent debate [79] to aggregate the reasoning capabilities of multiple LLMs. In this setup, gpt-3.5-turbo and GLM4 serve as the affirmative debater and negative debater, respectively, while gpt-4o-mini acts as the judge. We then analyze the same 50 cases mentioned above, with the results presented in the second row of Table IX. It can be observed that errors in intermediate reasoning steps are significantly reduced under this multi-LLM framework. Subsequently, we use the improved rationale to train the SLM, and its performance is reported in Table IX. As we can see, the improvement in rationale quality leads to an increase in SLM performance. Moreover, a comparison of Fig. 6 illustrates a reduction in the influence of LLM bias on SLM.
Besides, a filtering strategy based on answer alignment is widely used to avoid the influence of low-quality teacher outputs. Specifically, one can employ regex matching for strings to compare the conclusion of rationale generated by the teacher LLM against the ground-truth answer, retaining only those samples where the answers align. Since correct conclusions correspond to the high-quality rationale, this filtering strategy significantly enhances the quality of the rationale used for distillation, thereby reducing the adverse effects of suboptimal teacher performance. Although this strategy may reduce the amount of available data, prior studies (e.g., Step-by-Step [7]) have demonstrated that CoT distillation remains effective even when only a subset of samples with high-quality rationale is preserved. This is because high-quality rationale enables the student model to get answers based on logical reasoning, rather than superficial pattern matching between questions and answers that rely on large datasets to learn. We validate that the proposed ATM exhibits this property, as shown in Table XII. Specifically, we simulated scenarios involving sub-performing teacher LLMs by varying the proportion of retained samples. The results indicate that even when only 50% of the training data is retained, the student model still outperforms standard fine-tuning.
In summary, bias in the teacher LLM will affect the student SLM. Introducing a better reasoner has been proven to be an effective way to mitigate this issue, and the data filtering strategy play a significant role when the rationale of LLM is noisy.
TABLE XII: Effect of different training data rates on SQA. The rate refers to the proportion of data to be retained, and the student model is GPT2-Large. The accuracy (%) is reported.
Standard Finetuning ATM (rate=100%) ATM (rate=75%) ATM (rate=50%) ATM (rate=25%) ATM (rate=12.5%)
63.31 65.06 64.33 64.63 62.73 60.26
IV-F2 Scalability of ATM
The scalability of ATM depends on whether the perception module can be effective across different applications and languages. In the perception module in ATM, query tokens pass through a cross-attention block, which makes query tokens interact with both the question and additional extracted features $a_{i}$ before being output. As a result, ATM can dynamically adjust the generation of soft prompt tokens based on the characteristics of the input question. This means that as long as the training data includes samples from different applications or languages, the perception module can learn to adjust soft prompt tokens accordingly. Consequently, the ATM can be extended to various applications and languages theoretically. To further validate the scalability of soft prompt tuning in ATM, we construct a comprehensive cross-lingual and cross-task reasoning dataset, consisting of CMATH [80] (Chinese mathematical reasoning), MA (English mathematical reasoning), SQA (English commonsense reasoning), LLC (English symbolic reasoning), and TSO (English logical reasoning). The dataset is split into training and validation sets with a 7:3 ratio. The results in Table X demonstrate that ATM consistently outperforms other approaches, proving its scalability.
IV-F3 Impact of features fed into the perception module
Whether incorporating a diverse set of explicit features in $a_{i}$ in addition to word count can enhance the learning capability of the perception module is displayed in Table XI. In the table, readability is the perplexity of the input question calculated by GPT2-Large, while MDD is the Mean Dependency Distance of the input question. Readability and MDD are input into the model in the form of “Perplexity: num” and “Mean Dependency Distance: num”, respectively.
The results in Table XI indicate that: a) Feeding question tokens into the perception module provides the most significant performance gain for SLM; b) As more explicit features are incorporated into $a_{i}$, the model’s performance gradually improves, but the observed improvements are relatively minor; c) The increase in model complexity resulting from these features is also minimal. These findings suggest that the perception module primarily relies on the implicit, comprehensive features of the question itself rather than $a_{i}$ for complexity estimation. If computational resources allow, the inclusion of carefully designed explicit features can offer measurable benefits. Conversely, in scenarios that prioritize model simplicity, this component can be entirely omitted without causing significant performance degradation.
@@ V Conclusion
VI ACKNOWLEDGMENT
This work was supported by the National Science Fund for Distinguished Young Scholars of China (No. 62325604), and the National Natural Science Foundation of China (No. 62276271 and No. 62421002).